OpenStudy (anonymous):

factor if possible 24x^3y+70x^2y+50xy

OpenStudy (anonymous):

first factor out a 2xy

OpenStudy (anonymous):

24x^3y+70x^2y+50xy = 2xy[12x^2 + 35x + 25] =2xy[(4x + 5)(3x + 5)]

OpenStudy (anonymous):

taking common 2xy then it becomes easy

OpenStudy (anonymous):

how did you guys come out with that ....my confuson is when 2xy and 12x^2 are multiplied doesnt it come out to 24xy^3,,, so thee y goes after the exponent to the tird and not befor

OpenStudy (anonymous):

in multiply a*b=b*a so don't worry